OpenStudy (anonymous):

1B-2 c) Maybe it's because I've been at this all day, but I cannot figure out why s(b/32)= b^2/64 For those who do not wish to search for reference, b is initial velocity in f/s. s is distance in feet t is time in seconds s= bt−16t^2 v= ds/dt at t -> v=b-32t as h->0 so max height is at t=b/32 (the time that v=0) I understand that s(b/32) should give me the answer, but I must be missing something in the algebra. Danke

OpenStudy (anonymous):

s(b/32)=b*(b/32)-16*(b/32)^2 =b^2/32-b^2/64 s(b/32)=b^2/64... what's the problem??

OpenStudy (anonymous):

I see it now, had to sleep on it I guess. I finally broke everything down as powers of 2 and it was easy to see. Thank you.